OpenStudy (anonymous):

Find all the zeros of the function. f(x)=x^4+4x^3-21x^2

OpenStudy (kc_kennylau):

using synthetic division? @KellyN

OpenStudy (luigi0210):

Factor

OpenStudy (anonymous):

\(f(x) = x^4+4x^3-21x^2 = x^2\left(x^2+4x-21\right)\). Now set this equal to 0. \(x^2\left(x^2+4x-21\right) = 0\). So \(x^2 = 0 \) and \(x^2-4x-21 = 0\). \(x^2 = 0\) -> x = 0 multiplicity 2 \(x^2-4x-21 = 0\) -> \(\left(x+7\right)\left(x-3\right) = 0\) -> x = -7, x = 3. So the zeros of f(x) are 0, 0, -7, and 3.